Two cars are driving toward each other on a straight and level road. One car is traveling at 100 km/h north and the other car is traveling at 42.0 km/h south, both velocities measured relative to the road. At a certain instant, the distance between the cars is 10.0 km. How long will take for the two cars to meet?



1
Expert's answer
2021-09-18T14:58:43-0400

The distance traveled by the first car "x=v_1 \times t"

The distance traveled by the second car "d-x = v_2 \times t"

"v_1 \times t = d -v_2 \times t \\

d = v_1 \times t + v_2 \times t \\

d = t(v_1+v_2) \\

t = \frac{d}{v_1+v_2} \\

t = \frac{10.0}{100 + 42.0} = 0.0704 \;h \\

t = 0.0704 \times 60 \times 60 = 253 \;sec"